OpenStudy (anonymous):

Solve V = π r2h for h.

OpenStudy (anonymous):

V = 2(pi)rh Divide by 2(pi)r on both sides V/2(pi)r = 2(pi)rh/2(pi)r V/2(pi)r = h h = V/2(pi)r

OpenStudy (anonymous):

does that help you out

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

your trying to find H correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what it is telling you to do is divide both sides of the equation by 2(pi)r to get the value of H

OpenStudy (anonymous):

Look how the equation is written and it explains how it is done

OpenStudy (anonymous):

okay I get what you did now

OpenStudy (anonymous):

that is how for these equations if you are trying to find the value the correct way should be how it is done!

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

I have another question. Solve P = 2(l + w) for l. What are the missing values in the table? P w l 14 2 ? 22 8 ?

OpenStudy (anonymous):

okay here goes and follow these is best I can help you

OpenStudy (anonymous):

Basically distribute the 2 with the l and w, so you end up with an equation of P = 2l+2w. Since you have the values for P and W, you can just plug them in, so the only term left would be "l". Does that make it more clear?

OpenStudy (anonymous):

all you have to do is move the constants to one side and make "l" by itself, that would be your solution to "l".

OpenStudy (anonymous):

you had me up until you said plug in

OpenStudy (anonymous):

place the number that is already given

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

yeah I got p-2w/2=l but how do I plug them in

OpenStudy (anonymous):

type them in

OpenStudy (anonymous):

So 8 and 14 is my answer for l